# Exploration of morphological coherence in open clusters with “core-shell” structure

### v1

@@ 3.5 Special sample clusters
Figure 11 displays the distribution between the ERs of the sample clusters and their outer ellipticities. All colored dots (yellow and red) and different kinds of symbols (marked green and blue) represent sample clusters. We can see from Fig. 11 that most sample clusters (yellow dots) are located in 0.25 $\leq$ ER $\leq$ 1. However, several special clusters can be found in ER $<$ 0.25. Their external morphology is more elliptical than their internal morphology, which is more elliptical than most sample clusters (yellow dots). Among these sample clusters, there are seven sample clusters with ER less than 0.25 and its error bar extends to out of 0.25. Only three open clusters have ERs less than 0.25 with their error bars in the 0.25 range. They are marked by different kinds of symbols, which are NGC 1647 (green open pentagram), NGC 6949 (green open pentagon), and NGC 3532 (green open triangle), respectively, as shown in Fig. 11.
@@
Figure 12 shows the 3D spatial distribution of eight special clusters of our sample, as well as the fitted ellipsoids of their internal and external structures. From the upper left panel of this picture, we can see the 3D spatial structure of NGC 1647, including its internal and external structures. NGC 1647 is an intermediate-age open cluster (about 117 Myr), with 599 members with a probability more than or equal to 0.5 and 383 stars with a probability less than 0.5, and about 577 pc (Hunt & Reffert, 2023). Previous research (e.g., Hassan, 1972; Pismis & Bozkurt, 1977; Geffert et al., 1996; Zdanavičius et al., 2005; Guerrero et al., 2011; Sánchez et al., 2018) focused mainly on its photometry and proper motion, and the study of its morphology was rarely covered. Zdanavičius et al. (2005, see their Fig 7) reported NGC 1647 has a loose structure and is elongated in the direction roughly perpendicular to our galaxy, which was based on the star counts method in the infrared $K$ passband. However, in this work, we found that it has a compact internal structure and loose external morphology, with a small ER. This means that its external region is more elliptical than its internal area. Based on the 3D spatial distribution of its external region, we speculated that it may be undergoing stellar vaporization, with escaped stars gradually becoming members of the tidal tail.
@@
We can see from the upper median panel of Fig. 12 the 3D spatial structure of NGC 3532. It is an extremely rich cluster, located in the Carina field of the Milky Way (Eggen, 1981), with an intermediate age (about 238 Myr) and about 472 pc, has 1843 members with probability more than or equal to 0.5 and 1571 stars with probability less than 0.5 (Hunt & Reffert, 2023). Gieseking (1981) pointed out that NGC 3532 has an extended cluster halo. In the study, it presents a spherical internal structure and a slightly elliptical external system. Meanwhile, its small MD implies a high degree of morphological stability. Due to its dense internal core, there is a high probability of two-body interaction, which allows low-mass members to transfer into the outer regions of the cluster and causes massive stars to sink into its core area. Therefore, this leads to a loose external structure of NGC 3532.
@@
NGC 6494 in our sample is also a cluster with a more elliptical external shape than internal morphology, as shown in the upper right panel of Fig. 12. It, about 713 pc and an intermediate-age (about 226 Myr) (Hunt & Reffert, 2023), is an intermediate richness (Sanders & Schroeder, 1980), with 465 members with membership less than 0.5 and 663 stars with membership more than or equal to 0.5 (Hunt & Reffert, 2023). Similar to the 3D spatial structures of NGC 1647 and NGC 3532, NGC 6494 has also a dense internal region and a relatively loose external structure, with a small MD.
@@
Furthermore, we can also see in Fig. 11 several sample clusters that are opposite to the above clusters. They are Alessi 8 (blue open pentagram), CWNU 1205 (blue open triangle), OCSN 98 (blue open circle), HSC 381 (blue open pentagon), and Stock 23 (blue open inverted triangle), with their ERs greater than 1, while the error bars of their ER jump out of that range. Therefore, we try here to analyze the reasons for the instability of their ERs as well as the morphological stability of these clusters.
@@
OCSN 98, discovered firstly by Qin et al. (2023), is a close ($\sim$ 151 pc) and very young ($\sim$ 4.5 Myr) open cluster, with 75 members more than or equal to 0.5 and without members less than 0.5 (Hunt & Reffert, 2023). Interestingly, we found that it has a circular external structure and a slightly elliptical internal morphology, with a small MD, as displayed in the lower left panel of Fig. 12. Due to its loose external system, the ellipsoid fitting error of its external region is relatively large, thus causing also its large error in ER. Furthermore, It is strange for OCSN 98 to present the circular external morphology. Because of its low height (||||Z|||| $<$ 100 pc) and its small number of members, it gets a high probability of being stretched by the gravitational perturbation of the Galactic disk.
@@ 4 Summary
This work is based on the biggest census catalog of open clusters and their member star catalogs from Gaia-DR3, from which 132 open clusters with a 3D core-shell structure that fit the ellipsoidal model have been selected as our sample. Using sample clusters’ MDs and ERs to characterize their morphological coherence, we investigated the distribution of the morphological coherence (MD and ER) of the sample clusters, exploring them in relation to foundational parameters (age and members’ number), spatial positions, and an external environment. Finally, we analyzed the special samples with a small ER or a large ER. Some of the conclusions drawn from this work are as follows:
@@
For the first time, we investigated the morphological coherence (MD and ER) of an open cluster with a 3D core-shell structure. What we can be sure of is that the morphological coherence of open clusters would certainly change in the struggle between their gravitational binding and the constant perturbations of the external environment. In future studies of cluster morphology, it will become a trend to unravel the intertwined external environmental perturbations that affect the morphology of open clusters, as this will contribute to a clear understanding of the morphological evolution of open clusters. We acknowledged that after the submission of this work, a revised classification for the cluster catalog of Hunt & Reffert (2023) was published in Hunt & Reffert (2024), which may impact our samples and results. In addition, we also need to acknowledge that some sample clusters still have members in tidal tails, thus our fitting method is also adopted to these sample clusters with gravitationally bound and non-gravitational bound member stars. In the analysis of results, we did not distinguish between sample clusters with or without tidal tail members. In future research, we will explore these two types of open clusters separately.
@@ Acknowledgements.
A.1 The Ellipsoid fitting method
@@
The ellipsoid fitting method adopted in this study referred to the study of Li & Griffiths (2004). The main idea is to use the least squares ellipsoid fitting to calculate the optimal fitting ellipsoid by matrix operations. The equation involved is Eq. 4, which is the general ellipsoid equation. We substitute the data of a cluster to be fitted into this equation and obtain the coefficients of the general expression for the best-fitting ellipsoid (A, B, C, D, E, F, G, H, I) by matrix operations. The coefficients can then be used to calculate the three axes (a,b,c) in Eq. 5, which is a standard ellipsoid expression.
$Ax^{2}+By^{2}+Cz^{2}+2Dxy+2Exz+2Fyz+2Gx+2Hy+2Iz=1$ (4)
$\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}=1$ (5)
@@
With the ellipsoid fitting method, we can obtain the optimal fitting ellipsoid for any cluster to be fitted. But whether the optimal ellipsoid is a valid ellipsoid or not, we need to judge by the errors of the three axes. In this study, we calculate the errors ($\sigma_{a}$, $\sigma_{b}$, and $\sigma_{c}$) of the three axes of the fitting ellipsoid by the following Eqs. 6, 7, 8, 9, 10, 11, and 12. In order to make this calculation simple, we firstly transformed Eq. 5 into the format of Eq. 9 by using three equivalent substitutions (Eqs. 6, 7, and 8). Because Eq. 10 is the fitting error ($\sigma$) equation of Eq. 9, then it can be used to derive the error ($\sigma_{k}$) of $k$ in Eq. 9, see Eq. 11. Finally, we can obtain Eq. 12 to calculate the error of the $a$ axis of fitting ellipsoids by combining Eqs. 8 and 11. Similarly, we also can calculate the errors of the $b$ and $c$ axes based on the above process. Here, $x{}^{\prime}$, $y{}^{\prime}$, and $k$ are equivalent substitutes for physical quantities (3D spatial coordinates: $x$, $y$, and $z$, and three axes: $a$, $b$, and $c$), with $i$ and $N$ denoting members of clusters fitted and the number of the clusters’ members. We take the triaxial errors of the fitted ellipsoid less than the triaxial lengths, respectively, as a condition for valid ellipsoid fitting, which is also our criterion for the final determination of sample clusters.
$x{}^{\prime}=-x^{2}$ (6)
$y{}^{\prime}=\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}$ (7)
$k=\frac{1}{a^{2}}$ (8)
$y{}^{\prime}=kx{}^{\prime}+1$ (9)
$\sigma=\sqrt{\frac{\sum(y_{i}^{{}^{\prime}}-kx_{i}^{{}^{\prime}}-1)^{2}}{N-2}}$ (10)
$\sigma_{k}=\sigma\sqrt{\frac{N}{N\sum(x_{i}^{{}^{\prime}})^{2}-(\sum x_{i}^{{}^{\prime}})^{2}}}$ (11)
$\sigma_{a}=\frac{a^{3}}{2}\sigma_{k}$ (12)
A.2 The distance correction method
@@
This work uses the Bayesian distance correction method, which was originally systematically provided by Bailer-Jones (2015); Carrera et al. (2019), and then continued to be developed and used by Pang et al. (2021), also adopted by other works (e.g., Ye et al., 2021; Hu et al., 2023; Qin et al., 2023). The following formula mainly refers to the appendix of Carrera et al. (2019). Equation 13 is the posterior distribution function of the Bayesian distance correction model, while Eqs. 14 and  15 are the likelihood and prior distribution functions of this model, respectively.
$f(d\mid\omega)\propto f(\omega\mid d)\Pi(d)$ (13)
$f(\omega\mid d)\propto e^{\frac{-(\omega-\frac{1}{d})^{2}}{2\sigma_{\omega}^{2}}}$ (14)
$\Pi(d)=(1-\alpha)d^{2}e^{-\frac{d}{L}}+\frac{\alpha}{\sqrt{2\pi\sigma_{d}^{2}}}e^{-\frac{(d-d_{0})^{2}}{2\sigma_{d}^{2}}}$ (15)
@@
The introduction of important parameters in the above formulation is as follows. $\omega$ indicates each member parallax measurement, with $\sigma_{\omega}$ equal to the error bar of each member parallax measurement. $d_{0}$ denotes each cluster center, with L = 8 kpc (Bailer-Jones, 2015). $\alpha$ in the study represents the member probability. $\sigma_{d}$ is the standard deviation of the distances of all member stars of a cluster to the cluster center, which was also uesd by Pang et al. (2021), with $d$ being the corrected distance of each star. We refer the reader to Carrera et al. (2019); Pang et al. (2021) for a more detailed Bayesian distance correction method.
@@
In order to verify that our Bayesian approach is correct, we tested and validated our approach with the help of previous cluster data. Since there is no published corrected distance data for cluster member stars in Carrera et al. (2019), we carried out this test using the corrected distance data for cluster member stars in Pang et al. (2021). We downloaded and got the raw data of the open cluster NGC 2547 and the corrected distance of the cluster from the data published in Pang et al. (2021). We then tested the raw data of the cluster with our distance correction method and compared the obtained results with the corrected distances in Pang et al. (2021), as shown in Fig. 13. We can see that the results of our test are consistent with the results of Pang et al. (2021). Hence, our Bayesian distance correction approach is correct and effective.

### v2

@@ 3.5 Special sample clusters
Figure 11 displays the distribution between the ERs of the sample clusters and their outer ellipticities. All colored dots (yellow and red) and different kinds of symbols (marked green and blue) represent sample clusters. We can see from Fig. 11 that most sample clusters (yellow dots) are located in 0.25 $\leq$ ER $\leq$ 1. However, several special clusters can be found in ER $<$ 0.25. Their external morphology is more elliptical than their internal morphology, which is more elliptical than most sample clusters (yellow dots). Among these sample clusters, there are seven sample clusters with ER values lower than 0.25 and its error bar extends to out of 0.25. Only three open clusters have ERs less than 0.25 with their error bars in the 0.25 range. They are marked by different kinds of symbols, which are NGC 1647 (green open pentagram), NGC 6949 (green open pentagon), and NGC 3532 (green open triangle), respectively, as shown in Fig. 11.
@@
Figure 12 shows the 3D spatial distribution of eight special clusters of our sample, as well as the fitted ellipsoids of their internal and external structures. From the upper left panel of this picture, we can see the 3D spatial structure of NGC 1647, including its internal and external structures. NGC 1647 is an intermediate-age open cluster (about 117 Myr), with 599 members with a probability more than or equal to 0.5 and 383 stars with a probability less than 0.5, and about 577 pc (Hunt & Reffert, 2023). Previous research (e.g., Hassan, 1972; Pismis & Bozkurt, 1977; Geffert et al., 1996; Zdanavičius et al., 2005; Guerrero et al., 2011; Sánchez et al., 2018) has focused mainly on its photometry and proper motion, whereas the study of its morphology was rarely covered. Zdanavičius et al. (2005, see their Fig 7) reported NGC 1647 having a loose structure and elongated in the direction roughly perpendicular to our galaxy, which was based on the star counts method in the infrared $K$ passband. However, in this work, we found that it has a compact internal structure and loose external morphology, with a small ER. This means that its external region is more elliptical than its internal area. Based on the 3D spatial distribution of its external region, we speculated that it may be undergoing stellar vaporization, with escaped stars gradually becoming members of the tidal tail.
@@
We can see from the upper median panel of Fig. 12, the 3D spatial structure of NGC 3532. It is an extremely rich cluster, located in the Carina field of the Milky Way (Eggen, 1981), with an intermediate age (about 238 Myr) and about 472 pc, has 1843 members with probability more than or equal to 0.5 and 1571 stars with probability less than 0.5 (Hunt & Reffert, 2023). Gieseking (1981) pointed out that NGC 3532 has an extended cluster halo. In the study, it presents a spherical internal structure and a slightly elliptical external system. Meanwhile, its small MD implies a high degree of morphological stability. Due to its dense internal core, there is a high probability of two-body interaction, which allows low-mass members to transfer into the outer regions of the cluster and causes massive stars to sink into its core area. Therefore, this leads to a loose external structure of NGC 3532.
@@
NGC 6494 in our sample is also a cluster with a more elliptical external shape than internal morphology, as shown in the upper right panel of Fig. 12. At about 713 pc and an intermediate-age (about 226 Myr) (Hunt & Reffert, 2023), it displays an intermediate richness (Sanders & Schroeder, 1980), with 465 members with membership less than 0.5 and 663 stars with membership more than or equal to 0.5 (Hunt & Reffert, 2023). Resembling the 3D spatial structures of NGC 1647 and NGC 3532, NGC 6494 has also a dense internal region and a relatively loose external structure, with a small MD.
@@
Furthermore, we can also see in Fig. 11 several sample clusters that are opposite to the above clusters. They are Alessi 8 (blue open pentagram), CWNU 1205 (blue open triangle), OCSN 98 (blue open circle), HSC 381 (blue open pentagon), and Stock 23 (blue open inverted triangle), with their ERs greater than 1, while the error bars of their ER jump out of that range. Therefore, we try to analyze the reasons for the instability of their ERs as well as the morphological stability of these clusters.
@@
OCSN 98, first discovered by Qin et al. (2023), is a close ($\sim$ 151 pc) and very young ($\sim$ 4.5 Myr) open cluster, with 75 members more than or equal to 0.5 and without members less than 0.5 (Hunt & Reffert, 2023). Interestingly, we found that it has a circular external structure and a slightly elliptical internal morphology, with a small MD, as displayed in the lower left panel of Fig. 12. Due to its loose external system, the ellipsoid fitting error of its external region is relatively large, thus causing also its large error in ER. Furthermore, it is strange for OCSN 98 to present the circular external morphology. Because of its low height (||||Z|||| $<$ 100 pc) and its small number of members, it gets a high probability of being stretched by the gravitational perturbation of the Galactic disk.
@@ 4 Summary
This work is based on the biggest census catalog of open clusters and their member star catalogs from Gaia-DR3, from which 132 open clusters with a 3D core-shell structure that fit the ellipsoidal model have been selected as our sample. Using the sample clusters’ MDs and ERs to characterize their morphological coherence, we investigated the distribution of the morphological coherence (MD and ER) of the sample clusters, exploring them in relation to foundational parameters (age and member numbers), spatial positions, and an external environment. Finally, we analyzed the special samples with a small ER or a large ER. Some of the conclusions drawn from this work are as follows:
@@
For the first time, we investigated the morphological coherence (MD and ER) of an open cluster with a 3D core-shell structure. What we can be sure of is that the morphological coherence of open clusters would certainly change in the struggle between their gravitational binding and the constant perturbations of the external environment. In future studies of cluster morphology, it will become commonplace to unravel the intertwined external environmental perturbations that affect the morphology of open clusters, as this will contribute to a clear understanding of the morphological evolution of open clusters. We acknowledg that after the submission of this work, a revised classification for the cluster catalog of Hunt & Reffert (2023) was published in Hunt & Reffert (2024), which may impact our samples and results. In addition, we need to acknowledge that some sample clusters still have members in tidal tails, thus our fitting method is also adopted to these sample clusters with gravitationally bound and non-gravitational bound member stars. In the analysis of results, we did not distinguish between sample clusters with or without tidal tail members. In future research, we will explore these two types of open clusters separately.
@@ Appendix A Supplementary material
A.1 Ellipsoid fitting method
@@
The ellipsoid fitting method adopted in this study referred to the study of Li & Griffiths (2004). The main idea is to use the least squares ellipsoid fitting to calculate the optimal fitting ellipsoid by matrix operations. The equation involved is Eq. 4, which is the general ellipsoid equation. We substitute the data of a cluster to be fitted into this equation and obtain the coefficients of the general expression for the best-fitting ellipsoid (A, B, C, D, E, F, G, H, I) via matrix operations. The coefficients can then be used to calculate the three axes (a,b,c) in Eq. 5, which is a standard ellipsoid expression:
$Ax^{2}+By^{2}+Cz^{2}+2Dxy+2Exz+2Fyz+2Gx+2Hy+2Iz=1,$ (4)
$\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}}=1.$ (5)
@@
With the ellipsoid fitting method, we can obtain the optimal fitting ellipsoid for any cluster to be fitted. However, the question of whether the optimal ellipsoid is a valid ellipsoid or not needs to be judged by the errors of the three axes. In this study, we calculated the errors ($\sigma_{a}$, $\sigma_{b}$, and $\sigma_{c}$) of the three axes of the fitting ellipsoid by the following Eqs. 6, 7, 8, 9, 10, 11, and 12. In order to make this calculation simple, we firstly transformed Eq. 5 into the format of Eq. 9 by using three equivalent substitutions (Eqs. 6, 7, and 8). Because Eq. 10 is the fitting error ($\sigma$) equation of Eq. 9, then it can be used to derive the error ($\sigma_{k}$) of $k$ in Eq. 9, see Eq. 11. Finally, we can obtain Eq. 12 to calculate the error of the $a$ axis of fitting ellipsoids by combining Eqs. 8 and 11. Similarly, we also can calculate the errors of the $b$ and $c$ axes based on the above process. Here, $x{}^{\prime}$, $y{}^{\prime}$, and $k$ are equivalent substitutes for physical quantities (3D spatial coordinates: $x$, $y$, and $z$, and three axes: $a$, $b$, and $c$), with $i$ and $N$ denoting members of clusters fitted and the number of the clusters’ members. We take the triaxial errors of the fitted ellipsoid less than the triaxial lengths, respectively, as a condition for valid ellipsoid fitting, which is also our criterion for the final determination of sample clusters:
$x{}^{\prime}=-x^{2},$ (6)
$y{}^{\prime}=\frac{y^{2}}{b^{2}}+\frac{z^{2}}{c^{2}},$ (7)
$k=\frac{1}{a^{2}},$ (8)
$y{}^{\prime}=kx{}^{\prime}+1,$ (9)
$\sigma=\sqrt{\frac{\sum(y_{i}^{{}^{\prime}}-kx_{i}^{{}^{\prime}}-1)^{2}}{N-2}},$ (10)
$\sigma_{k}=\sigma\sqrt{\frac{N}{N\sum(x_{i}^{{}^{\prime}})^{2}-(\sum x_{i}^{{}^{\prime}})^{2}}},$ (11)
$\sigma_{a}=\frac{a^{3}}{2}\sigma_{k}.$ (12)
A.2 Distance correction method
@@
This work uses the Bayesian distance correction method, which was originally systematically provided by Bailer-Jones (2015) and Carrera et al. (2019), then continued to be developed and used by Pang et al. (2021) and also adopted by other works (e.g., Ye et al., 2021; Hu et al., 2023; Qin et al., 2023). The following formula mainly refers to the appendix of Carrera et al. (2019). Equation 13 is the posterior distribution function of the Bayesian distance correction model, while Eqs. 14 and  15 are the likelihood and prior distribution functions of this model, respectively. The equations are as follows:
$f(d\mid\omega)\propto f(\omega\mid d)\Pi(d),$ (13)
$f(\omega\mid d)\propto e^{\frac{-(\omega-\frac{1}{d})^{2}}{2\sigma_{\omega}^{2}}},$ (14)
$\Pi(d)=(1-\alpha)d^{2}e^{-\frac{d}{L}}+\frac{\alpha}{\sqrt{2\pi\sigma_{d}^{2}}}e^{-\frac{(d-d_{0})^{2}}{2\sigma_{d}^{2}}}.$ (15)
@@
The introduction of important parameters in the above formulation is as follows: $\omega$ indicates each member parallax measurement, with $\sigma_{\omega}$ equal to the error bar of each member parallax measurement; $d_{0}$ denotes each cluster center, with L = 8 kpc (Bailer-Jones, 2015); $\alpha$ represents the member probability; and $\sigma_{d}$ is the standard deviation of the distances of all member stars of a cluster to the cluster center, which was also uesd by Pang et al. (2021), with $d$ being the corrected distance of each star. We refer to Carrera et al. (2019); Pang et al. (2021) for a more detailed Bayesian distance correction method.
@@
To verify that our Bayesian approach is correct, we tested and validated our approach with the help of previous cluster data. Since there is no published corrected distance data for cluster member stars in Carrera et al. (2019), we carried out this test using the corrected distance data for cluster member stars in Pang et al. (2021). We downloaded and obtained the raw data of the open cluster NGC 2547 and the corrected distance of the cluster from the data published in Pang et al. (2021). We then tested the raw data of the cluster with our distance correction method and compared the obtained results with the corrected distances in Pang et al. (2021), as shown in Fig. 13. We can see that the results of our test are consistent with the results of Pang et al. (2021). Hence, our Bayesian distance correction approach is correct and effective.